# SAGA: Synthetic Audit Log Generation for APT Campaigns

### v1

@@ IV-C APT Intrusion Detection
APT intrusion detection is defined as the task of identifying potential APT attacks within an audit log, as established in previous studies [6, 7, 33]. To demonstrate the usefulness of SAGA datasets in supporting APT intrusion detection, we present the detection results using their own evaluation schemes of various methods and then discuss the causes of performance difference, considering the following factors: test granularity, Technique used, attack scenario, and duration. Table VI presents the performance of the three intrusion detection systems, AirTag (event-based), Kairos (time window-based) and Unicorn (graph-based) across various SAGA datasets. More details can be found in Tables S5, S6, and S7 of Supplementary Material B. Performance data for E3-CADETS, E3-THEIA, and E3-ClearScore were copied directly from their respective original papers. Unicorn, which analyzes the entire provenance graph to detect anomalies, generally performed well in both precision and recall. AirTag, on the other hand, which identifies intrusions by analyzing individual events, struggled with high rates of false positives and false negatives. Kairos, by grouping events within time windows, tended to reduce both false positives and false negatives, resulting in relatively low precision but higher recall.
@@
IV-D Evaluation on Technique Hunting
Technique hunting is the task of identifying potential threats, specifically Techniques, as defined by the MITRE ATT&CK framework. SAGA follows Algorithm 2 to generate attack campaigns, where each event is labeled with its corresponding Technique identifier. This structured synthetic audit log allows detection methods to effectively perform Technique hunting. The approaches are used for Technique hunting; one is SFM and the other is using Sigma rules. The performance results of Technique hunting for each APT attack campaign are reported in Table VII that details are given in Table S9 in the Supplementary Material C. The results indicate significant performance distinctions between SFM and Sigma rules, regardless of the generation method, attack scenario, or duration. SFM is capable of recognizing attack patterns by learning from synthetic data, whereas Sigma rules are limited to detecting attack patterns that match pre-defined signatures. This underscores the importance of the generated datasets in facilitating the advancement of learning-based model methodologies.
@@ IV-D Evaluation on Technique Hunting
On the other hand, while Sigma rules perform efficiently, they are based on contributions from human experts and may not encompass all attack behaviors, such as T1055.002 Portable Executable Injection (PEI) and T1491 Defacement (DF2) identified in this empirical study. As a result, certain malicious behaviors went undetected in the APT campaigns C-6, C-8, G-11, and G-13.
@@ IV-E APT Campaign Detection Evaluation
APT campaign detection involves identifying a potential APT campaign embedded in an audit log, which is characterized by a series of coordinated attack patterns. In line with the MITRE ATT&CK framework, an APT campaign can be described as a sequence of Techniques. SFM leverages this structure to detect the most likely APT campaign by comparing the discovered Techniques ($G_{q}$) against a pool of known APT campaigns ($G_{c}$). In this empirical study, $G_{c}$ includes 8 known campaigns and 30 randomly generated APT campaigns. Figure 7 illustrates the ranking of each SFM-identified campaign over three different durations, 15 minutes, 1 hour, and 1 day. SFM successfully identified most campaigns within the top 3 rankings for single campaign groups (C and G), across all three time durations. In the 15-minute scenarios, 65% of the campaigns were ranked first. One reason for this success could be that the temporal causal relationships is established among the Techniques discovered by SFM, which effectively distinguishes the most probable campaign from others. Furthermore, SFM demonstrated resilience to discrepancies between the discovered ability graph $G_{q}$ and the campaign Techniques $G_{c}$. However, the M group, consisting of multiple campaigns in a single scenario, posed a greater challenge compared to the groups of single campaigns. In addition, the results further suggest that as the duration of audit log grows, the complexity of identifying APT campaigns also increases. In conclusion, this evaluation highlights the potential of SAGA in effectively supporting the detection of APT campaigns.
@@ IV-F A Case Study of Unseen APT Campaign
However, these results highlight the practical value of SAGA datasets for training detection models and provide a solid foundation for refining dataset generation strategies. They also underscore the need for further improvements in APT campaign detection models to enhance their capability and accuracy. However, these results highlight the practical value of the SAGA dataset in supporting practical cybersecurity applications and providing a solid foundation for future improvements to dataset generation strategies. They also underscore the need for further improvements in APT campaign detection models to enhance their capability and accuracy.
@@ V Related Work
PWNJUTSU [41] and LADEMU [43] considered adversaries’ TTPs, defining the operational semantics of Techniques as specifications for procedures. PWNJUTSU utilized a semi-automated approach where each attack step was crafted by professional human Red Teamers, providing a highly skilled yet partially manual campaign construction process. In contrast, LADEMU offered a proof-of-concept implementation using the red-team emulator Caldera and demonstrated the use of a single attack scenario, APT29. Our study introduces a workflow of campaign generation with event logs, enabling fully automated APT simulations based on TTPs specified in the MITRE ATT&CK framework. This approach eliminates the need for manual intervention, ensuring more consistent and scalable attack simulations.
@@
VI Conclusion
This study presents a generative APT campaign framework, SAGA, designed to synthesize synthetic audit logs for nearly any configuration. These logs support the training of various machine learning-based APT-related detection machine learning models and facilitate the evaluation of various APT-related detection methods. To demonstrate the usefulness of SAGA’s synthetic datasets, a collection of empirical studies were conducted across multiple established baselines for APT intrusion detection, Technique hunting, and APT campaign detection. Although each has specific strengths, they also have limitations. A widely accepted dataset, like the synthetic audit log generated by SAGA, could play a crucial role in improving the accuracy and robustness of these detection methods.
However, the quality of SAGA datasets depends on access to extensive, high-quality campaign knowledge and diverse Technique implementations. The current implementation of SAGA uses 169 abilities and 80 Techniques, with ongoing efforts to expand these numbers to enhance the usefulness of SAGA datasets and cover more sophisticated APT campaigns. In this work, the Windows operating system is used as the target environment; however, the concept can be readily applied to other platforms as well.

### v2

@@ IV-C APT Intrusion Detection
APT intrusion detection is defined as the task of identifying potential APT attacks within an audit log, as established in previous studies [6, 7, 34]. To demonstrate the usefulness of SAGA datasets in supporting APT intrusion detection, we present the detection results using their own evaluation schemes of various methods and then discuss the causes of performance difference, considering the following factors: test granularity, Technique used, attack scenario, and duration. Table VI presents the performance of the three intrusion detection systems, AirTag (event-based), Kairos (time window-based) and Unicorn (graph-based) across various SAGA datasets. More details can be found in Tables S5, S6, and S7 of Supplementary Material B. Performance data for E3-CADETS, E3-THEIA, and E3-ClearScore were copied directly from their respective original papers. Unicorn, which analyzes the entire provenance graph to detect anomalies, generally performed well in both precision and recall. AirTag, on the other hand, which identifies intrusions by analyzing individual events, struggled with high rates of false positives and false negatives. When exposed to SAGA’s broader technique coverage, multi-stage campaign structure, and temporal variation, AirTag, a one-class SVM anomaly detection model built on a customized BERT embedding, struggles to find a stable decision boundary across such diverse contexts. Kairos was designed with specific assumptions about event sequence timing, which may not align with the more flexible and richly structured scenarios generated by SAGA. By grouping events within time windows, Kairos tended to reduce both false positives and false negatives, resulting in relatively low precision but higher recall.
@@
IV-D Evaluation on Technique Hunting
To demonstrate the practical utility of fine-grained Technique labeling, Technique hunting is the task of identifying potential threats, specifically Techniques, as defined by the MITRE ATT&CK framework. SAGA follows Algorithm 2 to generate attack campaigns, where each event is labeled with its corresponding Technique identifier. This structured synthetic audit log allows detection methods to effectively perform Technique hunting. The approaches are used for Technique hunting; one is SFM and the other is using Sigma rules. The performance results of Technique hunting for each APT attack campaign are reported in Table VII that details are given in Table S9 in the Supplementary Material C. The results indicate significant performance distinctions between SFM and Sigma rules, regardless of the generation method, attack scenario, or duration. SFM is capable of recognizing attack patterns by learning from synthetic data, whereas Sigma rules are limited to detecting attack patterns that match pre-defined signatures. This underscores the importance of the generated datasets in facilitating the advancement of learning-based model methodologies.
@@ IV-D Evaluation on Technique Hunting
On the other hand, while Sigma rules perform efficiently, they are based on contributions from human experts and may not encompass all attack behaviors, such as T1055.002 Portable Executable Injection (PEI) and T1491 Defacement (DF2) identified in this empirical study. As a result, certain malicious behaviors went undetected in the APT campaigns C-6, C-8, G-11, and G-13. These findings address the limitations of manually crafted, signature-based detection methods and highlight the importance of datasets like those generated by SAGA, which provide comprehensive logs from red-team emulation campaigns. Such datasets support the development and evaluation of both rule-based detection systems and learning-based models in a realistic environment.
@@ IV-E APT Campaign Detection Evaluation
APT campaign detection involves identifying a potential APT campaign embedded in an audit log, which is characterized by a series of coordinated attack patterns. In line with the MITRE ATT&CK framework, an APT campaign can be described as a sequence of Techniques. SFM leverages this structure to detect the most likely APT campaign by comparing the discovered Techniques ($G_{q}$) against a pool of known APT campaigns ($G_{c}$). In this empirical study, $G_{c}$ includes 8 known campaigns and 30 randomly generated APT campaigns. Figure 7 illustrates the ranking of each SFM-identified campaign over three different durations, 15 minutes, 1 hour, and 1 day. SFM successfully identified most campaigns within the top 3 rankings across all three time durations: 74% for 15 minutes, 84% for 1 hour, and 79% for 1 day. Notably, 50% of the campaigns were ranked first in both 1 hour and 1 day scenarios. One reason for this success could be that the temporal causal relationships is established among the Techniques discovered by SFM, which effectively distinguishes the most probable campaign from others. Furthermore, SFM demonstrated resilience to discrepancies between the discovered ability graph $G_{q}$ and the campaign Techniques $G_{c}$. However, the M group, consisting of multiple campaigns in a single scenario, posed a greater challenge compared to the groups of single campaigns. In addition, the results further suggest that as the duration of audit log grows, the complexity of identifying APT campaigns also increases. In conclusion, this evaluation highlights the potential of SAGA in effectively supporting the detection of APT campaigns.
@@ IV-F A Case Study of Unseen APT Campaign
However, these results highlight the practical value of the SAGA dataset in supporting practical cybersecurity applications and providing a solid foundation for future improvements to dataset generation strategies. They also underscore the need for further improvements in APT campaign detection models to enhance their capability and accuracy.
@@ V Related Work
PWNJUTSU [43] and LADEMU [45] considered adversaries’ TTPs, defining the operational semantics of Techniques as specifications for procedures. PWNJUTSU utilized a semi-automated approach where each attack step was crafted by professional human Red Teamers, providing a highly skilled yet partially manual campaign construction process. In contrast, LADEMU offered a proof-of-concept (PoC) implementation using the red-team emulator Caldera and demonstrated the use of a single attack scenario, APT29. Our study introduces a workflow of campaign generation with event logs, enabling fully automated APT simulations based on TTPs specified in the MITRE ATT&CK framework. This approach eliminates the need for manual intervention, ensuring more consistent and scalable attack simulations. Table VIII contrasts SAGA with previous framework and highlights SAGA’s unique advantages, particularly its configurability (in terms of duration, techniques, and campaign design), fine-grained labeling, and high degree of automation.
@@
TABLE VIII: Comparisons with Synthetic APT Dataset Generation.
Log
Generation
DAPT2020
[42]
PWNJUTSU
[43]
Unraveled
[44]
LADEMU
[45]
SAGA
Duration 3 months - 6 weeks 1 hour Configurable
# Attackers 1 22 3 1 Configurable
Attack Stage 4 5 5 4 6
# Attacks 16 13 15 79
197
Label
Stage,
Attack
Behavior
Stage,
Technique
Activity,
Stage,
Defender
Response,
Signature
Tactic,
Technique
Stage,
Technique,
Ability,
Campaign
Alignment
to
MITRE
ATT&CK
N Y Y Y Y
Threat
Model
Human
Experts
Human
Experts
Employee
Behavior
Generation
Red team
emulator
Red team
emulator
Automation
VI Discussion and Conclusion
This study presents a generative APT campaign framework, SAGA, designed to synthesize synthetic audit logs for nearly any configuration. These logs support the training of various machine learning-based APT-related detection machine learning models and facilitate the evaluation of various APT-related detection methods. To demonstrate the usefulness of SAGA’s synthetic datasets, a collection of empirical studies were conducted across multiple established baselines for APT intrusion detection, Technique hunting, and APT campaign detection. In light of the observation of artifact leakage, distributional shift, benign syntheticity, and attack covertness in the synthetic datasets [11], we emphasize that the SAGA-generated datasets are not limited to single-campaign training or evaluation. They can be flexibly composed to include multiple campaigns, time spans, and benign activities, allowing researchers to design training and testing splits that better reflect real-world conditions and capture diverse, multi-stage attack behaviors. Furthermore, SAGA enables researchers to explore distributional shifts by generating stage-specific audit logs, supporting the training and evaluation of models on early-stage versus late-stage attack behaviors.
The authenticity of SAGA-generated logs is grounded in two factors: the faithful implementation of attack techniques and the construction of realistic attack sequences. SAGA relies on a red-team emulator (i.e., Caldera) to execute concrete implementations of ATT&CK techniques. The more faithfully the emulator reproduces adversary behaviors, the more realistic the resulting logs. For sequence construction, SAGA supports both expert-instructed campaigns, emulating real-world APT scenarios as defined by domain experts (similar to manual red-teaming), and sampled sequences, assembled from statistically plausible and logically coherent technique transitions. This design enables SAGA to generate scalable and scenario-driven audit logs, addressing the broader need for labeled and scenario-driven log data for learning-based APT detection research.
The quality of SAGA datasets depends on access to extensive, high-quality campaign knowledge and diverse Technique implementations. The current implementation of SAGA uses 169 abilities and 80 Techniques, with ongoing efforts to expand these numbers to enhance the usefulness of SAGA datasets and cover more sophisticated APT campaigns. While the ATT&CK framework is continuously evolving, SAGA is designed to be extensible. New techniques can be integrated by either updating Caldera with additional abilities as they become available, or manually constructing new attack templates through expert-guided red-team emulation. Furthermore, we also recognize the potential of advanced tools such as large language models (LLMs) used in [54] to assist in the automatic attack generation – a promising direction for future research. Given the inherent differences between synthetic and real-world data, we caution that experimental results obtained using SAGA should be interpreted carefully, as synthetic logs may not present in real environments. As highlighted in [11], careful evaluation is necessary to avoid overestimating generalization capabilities due to unintended signals in synthetic datasets. In this work, the Windows operating system is used as the target environment; however, the concept can be readily applied to other platforms as well.